# The single edge notch fracture test for viscoelastic elastomers

### v1

@@ A Newtonian fluid ($\mu_{1}=\mu_{2}=\nu_{2}=0$, $\nu_{1}=+\infty$, $\eta=\eta_{0}$);
The prescription (11)-(12) has the additional merit of being able to describe the viscoelastic behavior of a wide range of elastomers, which typically exhibit non-Gaussian elasticityTreloar (2005); Lopez-Pamies (2010), as well as a strongly nonlinear viscosity that depends both on the deformation and on the deformation rateDoi and Edwards (1998); Bergström and Boyce (1998); Amin et al. (2006); Kumar and Lopez-Pamies (2016); Jianyou et al. (2018); Ghosh et al. (2021); Chockalingam et al. (2021); Chen and Ravi-Chandar (2022); Ricker et al. (2023). In particular, the viscosity of elastomers typically increases with increasing deformation and decreases with increasing deformation rate, the latter behavior being commonly referred to as shear-thinning. The viscosity function $\eta({\bf C},{\bf C}^{v})=\widetilde{\eta}(I_{1}^{v},\mathcal{J}_{2}^{{\rm NEq}})$ in (10) can be suitably prescribed to describe both of these behaviors via its dependence on deformation through $I_{1}^{v}$ and on deformation rate through $\mathcal{J}_{2}^{{\rm NEq}}$. In the parametric study that is presented next, we will use the specific constitutive prescriptionKumar and Lopez-Pamies (2016); Ghosh et al. (2021)
@@ 2.3 Initial and boundary conditions
$\left\{\begin{array}[]{l}{\bf y}({\bf X},0)={\bf X\vspace{0.2cm}}\\ p({\bf X},0)=\mu^{{\rm Eq}}+\mu^{{\rm NEq}}\vspace{0.2cm}\\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right.,\quad{\bf X}\in\overline{\mathrm{\Omega}}_{0},$
@@
[…]
@@
[…]
of the specimens are held firmly by stiff grips that are pulled apart in the ±$\pm{\bf e}_{3}$ directions resulting in a separation $l(t)$ between the grips. Consistent with the manner in which the vast majority of experiments are carried out, we restrict attention to the case when the grips are pulled apart at a constant rate $\dot{l}_{0}$, that is, when the separation between the grips is given by $l(t)=L+\dot{l}_{0}t$. This results in tests carried out at the constant global stretch rate $\dot{\mathrm{\Lambda}}_{0}=\dot{l}_{0}/L$. Therefore, making use of the notation $\partial\mathrm{\Omega}^{\mathcal{N}}_{0}=\partial\mathrm{\Omega}_{0}\setminus\left(\partial\mathrm{\Omega}^{\mathcal{B}}_{0}\cup\partial\mathrm{\Omega}^{\mathcal{T}}_{0}\right)$ and $\textbf{s}({\bf X},t)={\bf S}{\bf N}$, where ${\bf N}$ stands for the outward unit normal to the boundary $\partial\mathrm{\Omega}_{0}$, we have the boundary conditions
[…]
@@
$\left\{\begin{array}[]{l}\overline{y}^{\mathcal{B}}_{1}({\bf X},t)=X_{1}\vspace{0.15cm}\\ \overline{y}^{\mathcal{B}}_{2}({\bf X},t)=X_{2}\vspace{0.15cm}\\ \overline{y}^{\mathcal{B}}_{3}({\bf X},t)=X_{3}-\dfrac{\dot{l}_{0}}{2}t\end{array}\right.\quad{\rm and}\quad\left\{\begin{array}[]{l}\overline{y}^{\mathcal{T}}_{1}({\bf X},t)=X_{1}\vspace{0.15cm}\\ \overline{y}^{\mathcal{T}}_{2}({\bf X},t)=X_{2}\vspace{0.15cm}\\ \overline{y}^{\mathcal{T}}_{3}({\bf X},t)=X_{3}+\dfrac{\dot{l}_{0}}{2}t\end{array}\right..$
@@ 2.4 Governing equations
Neglecting inertia and body forces, the equations of balance of linear and angular momenta read as ${\rm Div}\,{\bf S}={\bf 0}$ and ${\bf S}{\bf F}^{T}={\bf F}{\bf S}^{T}$ for $({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]$. The latter is automatically satisfiedKumar and Lopez-Pamies (2016); Sadik and Yavari (2024) by virtue of the objectivity of the thermodynamic potentials (6) and (7) and so, upon combining all the above ingredients together and introducing the notation
${\bf S}({\bf X},t)=2\psi^{{\rm Eq}}_{I_{1}}\nabla{\bf y}+2\psi^{{\rm NEq}}_{I^{e}_{1}}\nabla{\bf y}{{\bf C}^{v}}^{-1}-p\nabla{\bf y}^{-T}$
@@
$\tau({\bf X},t)=\dfrac{\widetilde{\eta}(I_{1}^{v},\mathcal{J}_{2}^{{\rm NEq}})}{2\psi^{{\rm NEq}}_{I^{e}_{1}}}$
@@
$\left\{\begin{array}[]{ll}{\rm Div}\,{\bf S}={\bf 0},&({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]\\ \\ \det\nabla{\bf y}=1,&({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]\\ \\ {\bf y}({\bf X},t)=\overline{{\bf y}}^{\mathcal{B}}({\bf X},t),&({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{B}}_{0}\times[0,T]\\ \\ {\bf y}({\bf X},t)=\overline{{\bf y}}^{\mathcal{T}}({\bf X},t),&({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{T}}_{0}\times[0,T]\\ \\ {\bf S}{\bf N}=\textbf{0},&({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{N}}_{0}\times[0,T]\\ \\ {\bf y}({\bf X},0)={\bf X},&{\bf X}\in\mathrm{\Omega}_{0}\\ \\ p({\bf X},0)=\mu^{{\rm Eq}}+\mu^{{\rm NEq}},&{\bf X}\in\mathrm{\Omega}_{0}\end{array}\right.$ (14)
@@
$\left\{\begin{array}[]{ll}\dot{{\bf C}}^{v}({\bf X},t)=\dfrac{1}{\tau}\left(\nabla{\bf y}^{T}\nabla{\bf y}-\dfrac{1}{3}\left(\nabla{\bf y}^{T}\nabla{\bf y}\cdot{{\bf C}^{v}}^{-1}\right){\bf C}^{v}\right),&\\ \\ &({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]\\ \\ {\bf C}^{v}({\bf X},0)={\bf I},&{\bf X}\in\mathrm{\Omega}_{0}\end{array}\right.$ (15)
@@
All the results that we present below are generated by solving numerically the initial-boundary-value problem (14)-(15) via a variant of the scheme introduced by Ghosh et al.Ghosh et al. (2021), which is based on a Crouzeix-Raviart FE discretization of space and a high-order explicit Runge-Kutta finite difference discretization of time, as well as by a related scheme that has been recently implemented as an Abaqus UMAT subroutineLefèvre et al. ; we also generated corresponding 2D plane-stress results, which turned out to be in agreement with the fully 3D results. FEniCSx and Abaqus implementations of these numerical schemes are available on GitHub. [^1] http://pamies.cee.illinois.edu/Subroutines.html.
@@ 3 The Griffith criticality condition (2) for the single edge notch fracture test
    $-\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial{\mathrm{\Gamma}_{0}}}=-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda})$ (17)
@@
    $-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda}_{c})=G_{c}$ (18)
@@
$T_{c}=G_{c}-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{{\rm NEq}}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda}_{c})-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{v}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda}_{c})$
@@
$\mathcal{W}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda})=\displaystyle\int_{L}^{\mathrm{\Lambda}L}P\,{\rm d}l$
@@
By way of an example, Fig. 3 provides a representative result for the equilibrium stored-energy function $\psi^{{\rm Eq}}(I_{1})$ over the initial $\mathrm{\Omega}_{0}$ and the current $\mathrm{\Omega}(t)$ configurations of a specimen, with pre-existing crack of length $A=2$ mm, at a global stretch $\mathrm{\Lambda}=1.5$, that has been stretched at the global stretch rate $\dot{\mathrm{\Lambda}}_{0}=10^{0}$ s-1; see also Fig. S1 in the Supplementary Material. Figure 4 presents corresponding results for the equilibrium energy (16) and its derivative (17) plotted as functions of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ and the initial crack size $A$ at a fixed global stretch $\mathrm{\Lambda}$. Finally, Fig. 5 presents the derivative (17) of the equilibrium energy as a function of the global stretch $\mathrm{\Lambda}$ and the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ for a fixed initial crack size $A$. The results in Figs. 3 through 5 pertain to one of the single edge notch fracture tests considered in Section 4, that of an elastomer with material constants $\mu_{1}=0.09$ MPa, $\alpha_{1}=0.5$, $\mu_{2}=0.01$ MPa, $\alpha_{2}=2.5$, $\nu_{1}=1.8$ MPa, $\beta_{1}=0.5$, $\nu_{2}=0.2$ MPa, $\beta_{2}=7.5$, $\eta_{0}=N_{0}=\eta_{\infty}=2$ MPa s, and specimens of length $L=15$ mm, width $H=4$ mm, and thickness $B=0.5$ mm.
To illustrate the evaluation of the Griffith criticality condition (18), Fig. 5 includes a plane of constant value — in this case, $G_{c}=150$ N/m — whose intersection with the surface of the derivative $-\partial\mathcal{W}^{{\rm Eq}}/\partial{\mathrm{\Gamma}_{0}}$ allows to readily identify the critical global stretches $\mathrm{\Lambda}_{c}$ at which the crack grows for a given critical energy release rate $G_{c}$ as a function of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$.
@@
$A=1.9,2,2.1\;\rm{mm}.$
These specific values for $L,H,B,A$ are chosen here because they are representative of those used in experimentsHamed and Park (1999); Cristiano et al. (2011); Fan et al. (2017); Roucou et al. (2018); Slootman et al. (2020); Menn (2022); Wang and Fan (2023); Xue et al. (2024).
As also representative of the behavior of typical elastomersLopez-Pamies (2010); Breedlove et al. (2024), we make use of the material constants
@@
$\left\{\begin{array}[]{rl}\texttt{Weaker Growth:}&\left\{\begin{array}[]{l}\nu_{1}=0.9\nu,\;\beta_{1}=0.5\\ \nu_{2}=0.1\nu,\;\beta_{2}=1\end{array}\right.\vspace{0.18cm}\\ \texttt{Equal Growth:}&\left\{\begin{array}[]{l}\nu_{1}=0.9\nu,\;\beta_{1}=0.5\\ \nu_{2}=0.1\nu,\;\beta_{2}=2.5\end{array}\right.\vspace{0.18cm}\\ \texttt{Stronger Growth:}&\left\{\begin{array}[]{l}\nu_{1}=0.9\nu,\;\beta_{1}=0.5\\ \nu_{2}=0.1\nu,\;\beta_{2}=7.5\end{array}\right.\end{array}\right.$ (23)
@@
in the non-equilibrium stored-energy function (10)2. As labeled in (23), these material constants yield non-equilibrium stored-energy functions (10)2 with weaker ($\beta_{2}<\alpha_{2}$), equal ($\beta_{2}=\alpha_{2}$), and stronger ($\beta_{2}>\alpha_{2}$) growth conditions — in the limit of large deformations as $||{\bf F}||=\sqrt{I_{1}}\nearrow+\infty$ — than that of the equilibrium stored-energy function (10)1 with (22).
@@
$\left\{\begin{array}[]{rl}\texttt{Constant Viscosity:}&\left\{\begin{array}[]{l}N_{0}=\eta_{0}\\ \eta_{\infty}=\eta_{0}\end{array}\right.\vspace{0.18cm}\\ \texttt{Shear-thinning Viscosity:}&\left\{\begin{array}[]{l}N_{0}=\eta_{0}\\ \eta_{\infty}=0.01\eta_{0}\\ K_{2}=500\,{\rm MPa}^{-2},\;\gamma_{2}=1\end{array}\right.\vspace{0.18cm}\\ \texttt{Deformation-dependent}&\vspace{-0.3cm}\\ \texttt{Viscosity:}&\left\{\begin{array}[]{l}N_{0}=100\eta_{0}\\ K_{1}=2,\;\gamma_{1}=4\\ K_{2}=0\end{array}\right.\end{array}\right.$ (25)
@@
$\dot{\mathrm{\Lambda}}_{0}=0+,10^{-3},10^{-2},10^{-1},10^{0},10^{1},10^{2},10^{3},+\infty\;{\rm s}^{-1},$

### v4

@@ A Newtonian fluid ($\mu_{1}=\mu_{2}=\nu_{2}=0$, $\nu_{1}=+\infty$, $\eta=\eta_{0}$);
The prescription (24)-(25) has the additional merit of being able to describe the viscoelastic behavior of a wide range of elastomers, which typically exhibit non-Gaussian elasticity [Treloar2005, LP10], as well as a strongly nonlinear viscosity that depends both on the deformation and on the deformation rate [Doi98, BB98, Lion2006, KLP16, Khayat18, GSKLP21, Cohen21, Ravi22, Ricker23]. In particular, the viscosity of elastomers typically increases with increasing deformation and decreases with increasing deformation rate, the latter behavior being commonly referred to as shear-thinning. The viscosity function $\eta({\bf C},{\bf C}^{v})=\widetilde{\eta}(I_{1}^{v},\mathcal{J}_{2}^{{\rm NEq}})$ in (23) can be suitably prescribed to describe both of these behaviors via its dependence on deformation through $I_{1}^{v}$ and on deformation rate through $\mathcal{J}_{2}^{{\rm NEq}}$. In the parametric study that is presented next, we will use the specific constitutive prescription [KLP16, GSKLP21]
@@ 2.3 Initial and boundary conditions
$\left\{\begin{array}[]{l}{\bf y}({\bf X},0)={\bf X}\vskip 5.69046pt\\ p({\bf X},0)=\mu^{{\rm Eq}}+\mu^{{\rm NEq}}\vskip 5.69046pt\\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right.,\quad{\bf X}\in\overline{\mathrm{\Omega}}_{0},$ (27)
@@
$\partial\mathrm{\Omega}^{\mathcal{B}}_{0}=$ $\left\{{\bf X}:|X_{1}|\leq\dfrac{B}{2},\,0\leq X_{2}\leq H,\,X_{3}=-\dfrac{L}{2}\right\}$ (28)
@@
$\partial\mathrm{\Omega}^{\mathcal{T}}_{0}=$ $\left\{{\bf X}:|X_{1}|\leq\dfrac{B}{2},\,0\leq X_{2}\leq H,\,X_{3}=\dfrac{L}{2}\right\}$ (29)
of the specimens are held firmly by stiff grips that are pulled apart in the $\pm{\bf e}_{3}$ directions resulting in a separation $l(t)$ between the grips. Consistent with the manner in which the vast majority of experiments are carried out, we restrict attention to the case when the grips are pulled apart at a constant rate $\dot{l}_{0}$, that is, when the separation between the grips is given by $l(t)=L+\dot{l}_{0}t$. This results in tests carried out at the constant global stretch rate $\dot{\mathrm{\Lambda}}_{0}=\dot{l}_{0}/L$. Therefore, making use of the notation $\partial\mathrm{\Omega}^{\mathcal{N}}_{0}=\partial\mathrm{\Omega}_{0}\setminus\left(\partial\mathrm{\Omega}^{\mathcal{B}}_{0}\cup\partial\mathrm{\Omega}^{\mathcal{T}}_{0}\right)$ and $\textbf{s}({\bf X},t)={\bf S}{\bf N}$, where $\bf N$} stands for the outward unit normal to the boundary $\partial\mathrm{\Omega}_{0}$, we have the boundary conditions
$\left\{\begin{array}[]{ll}{\bf y}({\bf X},t)=\overline{{\bf y}}^{\mathcal{B}}({\bf X},t),&\hskip 4.26773pt({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{B}}_{0}\times[0,T]\vskip 4.26773pt\\ {\bf y}({\bf X},t)=\overline{{\bf y}}^{\mathcal{T}}({\bf X},t),&\hskip 4.26773pt({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{T}}_{0}\times[0,T]\vskip 4.26773pt\\ \textbf{s}({\bf X},t)={\bf 0},&\hskip 4.26773pt({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{N}}_{0}\times[0,T]\end{array}\right$. (30)
@@
$\left\{\begin{array}[]{l}\overline{y}^{\mathcal{B}}_{1}({\bf X},t)=X_{1}\vskip 4.26773pt\\ \overline{y}^{\mathcal{B}}_{2}({\bf X},t)=X_{2}\vskip 4.26773pt\\ \overline{y}^{\mathcal{B}}_{3}({\bf X},t)=X_{3}-\dfrac{\dot{l}_{0}}{2}t\end{array}\right.\quad{\rm and}\quad\left\{\begin{array}[]{l}\overline{y}^{\mathcal{T}}_{1}({\bf X},t)=X_{1}\vskip 4.26773pt\\ \overline{y}^{\mathcal{T}}_{2}({\bf X},t)=X_{2}\vskip 4.26773pt\\ \overline{y}^{\mathcal{T}}_{3}({\bf X},t)=X_{3}+\dfrac{\dot{l}_{0}}{2}t\end{array}\right.$. (31)
@@ 2.4 Governing equations
Neglecting inertia and body forces, the equations of balance of linear and angular momenta read as ${\rm Div}\,{\bf S}={\bf 0}$ and ${\bf S}{\bf F}^{T}={\bf F}{\bf S}^{T}$ for $({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]$. The latter is automatically satisfied [KLP16, Yavari24] by virtue of the objectivity of the thermodynamic potentials (14) and (15) and so, upon combining all the above ingredients together and introducing the notation
${\bf S}({\bf X},t)=2\psi^{{\rm Eq}}_{I_{1}}\nabla{\bf y}+2\psi^{{\rm NEq}}_{I^{e}_{1}}\nabla{\bf y}{{\bf C}^{v}}^{-1}-p\nabla{\bf y}^{-T}$ (32)
@@
$\tau({\bf X},t)=\dfrac{\widetilde{\eta}(I_{1}^{v},\mathcal{J}_{2}^{{\rm NEq}})}{2\psi^{{\rm NEq}}_{I^{e}_{1}}}$ (33)
@@
$\left\{\begin{array}[]{ll}{\rm Div}\,{\bf S}={\bf 0},&({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]\\ \det\nabla{\bf y}=1,&({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]\\ {\bf y}({\bf X},t)=\overline{{\bf y}}^{\mathcal{B}}({\bf X},t),&({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{B}}_{0}\times[0,T]\\ {\bf y}({\bf X},t)=\overline{{\bf y}}^{\mathcal{T}}({\bf X},t),&({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{T}}_{0}\times[0,T]\\ {\bf S}{\bf N}={\bf 0},&({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{N}}_{0}\times[0,T]\\ {\bf y}({\bf X},0)={\bf X},&{\bf X}\in\mathrm{\Omega}_{0}\\ p({\bf X},0)=\mu^{{\rm Eq}}+\mu^{{\rm NEq}},&{\bf X}\in\mathrm{\Omega}_{0}\end{array}\right$. (34)
@@
$\left\{\begin{array}[]{ll}\dot{{\bf C}}^{v}({\bf X},t)=\dfrac{1}{\tau}\left(\nabla{\bf y}^{T}\nabla{\bf y}-\dfrac{1}{3}\left(\nabla{\bf y}^{T}\nabla{\bf y}\cdot{{\bf C}^{v}}^{-1}\right){\bf C}^{v}\right),&({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]\\ {\bf C}^{v}({\bf X},0)={\bf I},&{\bf X}\in\mathrm{\Omega}_{0}\end{array}\right$. (35)
@@
All the results that we present below are generated by solving numerically the initial-boundary-value problem (34)-(35) via a variant of the scheme introduced by [GSKLP21], which is based on a Crouzeix-Raviart FE discretization of space and a high-order explicit Runge-Kutta finite difference discretization of time, as well as by a related scheme that has been recently implemented as an Abaqus UMAT subroutine [LSLP24]; we also generated corresponding 2D plane-stress results, which turned out to be in agreement with the fully 3D results. FEniCSx and Abaqus implementations of these numerical schemes are available on GitHub. [^1] http://pamies.cee.illinois.edu/repositories/ In all the simulations, we made use of FE meshes that were refined around the crack front with a minimum mesh size of $h=B/10$ and loading steps that were small enough to resolve the pertinent relaxation time of the elastomer; see, e.g., Remark 10 in [LSLP24].
@@ 3 The Griffith criticality condition (2) for the single edge notch fracture test
    $\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial{\mathrm{\Gamma}_{0}}}=-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda})$ (37)
@@
    $\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda}_{c})=G_{c}$ (38)
@@
Remark 3.2.
In practice, it suffices to make use of a finite-difference approximation to compute the derivative in (37). In this work, we make use of the central three-point finite-difference quotient
$\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial{A}}=\dfrac{\mathcal{W}^{{\rm Eq}}(\dot{\mathrm{\Lambda}}_{0},A+\Delta A;\mathrm{\Lambda})-\mathcal{W}^{{\rm Eq}}(\dot{\mathrm{\Lambda}}_{0},A-\Delta A;\mathrm{\Lambda})}{2\Delta A}$
with $\Delta A=A/20$, which was checked to generate converged results.
@@
$T_{c}=G_{c}-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{{\rm NEq}}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda}_{c})-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{v}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda}_{c})$ (39)
@@
$\mathcal{W}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda})=\displaystyle\int_{L}^{\mathrm{\Lambda}L}P\,{\rm d}l$ (43)
@@
By way of an example, Fig. 3 provides a representative result for the equilibrium stored-energy function $\psi^{{\rm Eq}}(I_{1})$ over the initial $\mathrm{\Omega}_{0}$ and the current $\mathrm{\Omega}(t)$ configurations of a specimen, with pre-existing crack of length $A=2$ mm, at a global stretch $\mathrm{\Lambda}=1.5$, that has been stretched at the global stretch rate $\dot{\mathrm{\Lambda}}_{0}=10^{0}$ s-1; see also Fig. 14 in the Appendix. Figure 4 presents corresponding results for the equilibrium energy (36) and its derivative (37) plotted as functions of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ and the initial crack size $A$ at a fixed global stretch $\mathrm{\Lambda}$. Finally, Fig. 5 presents the derivative (37) of the equilibrium energy as a function of the global stretch $\mathrm{\Lambda}$ and the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ for a fixed initial crack size $A$. The results in Figs. 3 through 5 pertain to one of the single edge notch fracture tests considered in Section 4, that of an elastomer with material constants $\mu_{1}=0.09$ MPa, $\alpha_{1}=0.5$, $\mu_{2}=0.01$ MPa, $\alpha_{2}=2.5$, $\nu_{1}=1.8$ MPa, $\beta_{1}=0.5$, $\nu_{2}=0.2$ MPa, $\beta_{2}=7.5$, $\eta_{0}=N_{0}=\eta_{\infty}=2$ MPa s, and specimens of length $L=15$ mm, width $H=4$ mm, and thickness $B=0.5$ mm.
To illustrate the evaluation of the Griffith criticality condition (38), Fig. 5 includes a plane of constant value — in this case, $G_{c}=150$ N/m — whose intersection with the surface of the derivative $\partial\mathcal{W}^{{\rm Eq}}/\partial{\mathrm{\Gamma}_{0}}$ allows to readily identify the critical global stretches $\mathrm{\Lambda}_{c}$ at which the crack grows for a given critical energy release rate $G_{c}$ as a function of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$.
@@
$A=1.9,2,2.1\;\rm{mm}.$ (44)
These specific values for $L,H,B,A$ are chosen here because they are representative of those used in experiments [Hamed99, Creton11, Cai17, Mbiakop18, Creton20, LeMenn22, Wang23, Dufresne24].
As also representative of the behavior of typical elastomers [LP10, BCLLP24], we make use of the material constants
@@
[…]
@@
in the non-equilibrium stored-energy function (23)2. As labeled in (46), these material constants yield non-equilibrium stored-energy functions (23)2 with weaker ($\beta_{2}<\alpha_{2}$), equal ($\beta_{2}=\alpha_{2}$), and stronger ($\beta_{2}>\alpha_{2}$) growth conditions — in the limit of large deformations as ${\bf F}||=\sqrt{I_{1}}\nearrow+\infty$ — than that of the equilibrium stored-energy function (23)1 with (45).
@@
[…]
@@
$\dot{\mathrm{\Lambda}}_{0}=0+,10^{-3},10^{-2},10^{-1},10^{0},10^{1},10^{2},10^{3},+\infty\;{\rm s}^{-1},$ (50)